Solve the following word problems. Be sure to label the unknowns and to show the equation you use for the solution.
Business and finance. The Randolphs used 12 more gal of fuel oil in October than in September and twice as much oil in November as in September. If they used 132 gal for the 3 months, ho much was used during each month?

Answer by stanbon(48558)

The Randolphs used 12 more gal of fuel oil in October than in September and twice as much oil in November as in September. If they used 132 gal for the 3 months, how much was used during each month?
--------------
Let Sept fuel use be x gallons.
October fuel use is x+12 gallons.
November fuel use is 2x gallons.
--------------
EQUATION:
x+x+12+2x=132
4x = 120
x=30 (Sept. use)
x+12=42 (Oct use)
2x=60 (Nov use)
=======================
Cheers,
Stan H.